OpenStudy (anonymous):

Average cost suppose that the daily cost C of manufacturing bicycles is given by C (x)=800x+5000. How many bicylces must be produce each day for the average cost to be more than $100.00

OpenStudy (johnweldon1993):

Are you sure it is not \[C(x) = 800x - 5000\]?

OpenStudy (anonymous):

yes it is, I just looking at that, I apologize for that...

OpenStudy (johnweldon1993):

Okay....well then it wants you to find how many bicycles need to be produced in order to turn a profit of at least $100 per day...so \[100 = 800x - 5000\] Can you solve that equation?

OpenStudy (anonymous):

Okay, so do I take 5000 and minus 800?

OpenStudy (johnweldon1993):

Notice how 800...has a variable 'x' with it....so you cannot do anything to that number right now.... your first step should be to isolate 800x on the right side...so try adding 5000 to both sides...you'll then have 5100 = 800x Now how would you solve for 'x'...?

OpenStudy (anonymous):

wow how did you get that? I can't solve this equation, sorry, not good @ word problems.... Solve for x??????/??

OpenStudy (johnweldon1993):

Don't worry about it.....here I'll show you

OpenStudy (johnweldon1993):

Or maybe @kropot72 can explain it better...

OpenStudy (anonymous):

but I am taking notes..... eager to learn

OpenStudy (anonymous):

no your doing just fine.... no matter what doing math without being in a classroom setting is very difficult to learn online.. Dealing with math needs to be hands on learning...but I am in college taking online classes, never knew I had to take math....

OpenStudy (kropot72):

The total cost function is normally expressed as \[C(x)=variable\ cost\ term+fixed\ cost\ term\] The question would make sense if the equation for total daily cost of production had been written as \[C(x)=80x+5000\] where x is the quantity produced in a day. @pizzaface Are you sure of the equation for your question?

OpenStudy (anonymous):

yes.... does it look wrong?

OpenStudy (anonymous):

okay it also has 80x+5000/x

OpenStudy (johnweldon1993):

Wait....the equation is \[C(x) = 80x + \frac{ 5000 }{ x }\]?

OpenStudy (anonymous):

no 80x+5000 divided by x

OpenStudy (kropot72):

The average cost of producing a bicycle is \[\frac{80x+5000}{x}\] @pizzaface Is that correct?

OpenStudy (anonymous):

yessssss:)

OpenStudy (kropot72):

So in that case the equation to solve to find the quantity x that would give an average cost per bicycle of exactly $100 is \[\frac{80x+5000}{x}=100...............(1)\] Multiplying both sides by x gives \[80x+5000=100x.........(2)\] Rearranging (2) we get \[20x=5000.........(3)\] If you divide both sides of equation (3) by 20 you will get a value for x. The number of bicycles produced per day (n) for the average cost to be more than $100 is n < x

OpenStudy (anonymous):

okay so is that my answer?

OpenStudy (kropot72):

We normally help questioners to arrive at answers by finding where the difficulties are and helping with explanations if we can. Are you able to divide both sides of the following equation by 20? 20x = 5000 If you successfully do that, you will almost have the answer.

OpenStudy (anonymous):

yes I think....

OpenStudy (kropot72):

Great! So what do you get?

OpenStudy (anonymous):

okay I divided 20 by both sides I got 250 (I hope that's right)

OpenStudy (kropot72):

Good :) So we get x = 250 bicycles for an average cost of $100.00 each. Therefore for the average cost to be more than $100.00, fewer than 250 bicycles would need to be produced.

OpenStudy (anonymous):

oh okay.. That's it wow, this was simple. now that know to figure out the formula... Thanx so much

OpenStudy (kropot72):

You're welcome :)